You want to paint the walls and ceiling of a room that measures 3mx4m and is 2.7m high . One liter tin of paint covers 8 square meters . How many tins of paint will you need ?

Answer by Boreal(15235)

4 walls
2 are 3*2.7=8.1 m^2*2=16.2 m^2
2 are 4*2.7=10.8 m^2*2=21.6 m^2
the ceiling is 3*4=12 m^2
The sum is 49.8 m^2
Six tins cover 48 m^2.
Seven tins are needed.